# The CoExplorer Technology Probe: A Generative AI-Powered Adaptive Interface to Support Intentionality in Planning and Running Video Meetings

### v1

@@ Abstract.
journalyear: 2024ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in interaction designccs: Human-centered computing Collaborative and social computing systems and tools
@@ 6.2.1. Linearity of meetings
Our findings suggest that both linear and non-linear meetings are common for our study participants, leading to mixed responses about surfacing the implicit phase structure in meetings. Some participants valued the structure that CoExplorer imposed on meetings by surfacing meeting goals and phases, helping to clarify attendee perceptions in advance of meetings, and supporting organizers in driving the meeting. This was perceived to improve meeting clarity and therefore productivity. However, some participants expressed concerns about the excessive structure imposed by CoExplorer, worrying that it may limit the natural spontaneity and flexibility of conversations. Considering these individual perceptions and preferences, it may be crucial for meeting support systems to capture the general consensus on the desired level of structure and adapt the workflow accordingly, rather than adopting a one-size-fits-all approach. Reconciling the variations in the group consensus and the individual preferences is a direction that needs further research. Systems could also help people decide on how much structure may be needed depending on their goals. It also points out the value in developing ways to adaptively adjust the level of structure during meetings. For instance, when people are moving through phases smoothly, the suggestion of structure could be more restrained, whereas if the meeting is being side-tracked, structure could be suggested in a stronger way by tracking the progress of the meeting through an intelligent system.
Design implication 2: The design of intelligent meeting systems should consider ways to address the tension between structured, linear meetings enabling intentionality and clarity, and less structured, non-linear meetings, enabling flexibility. This includes capturing attendees’ preferences about structure as well as helping decide how much structure may be beneficial for their goals. Variations in need for structure during a given meeting should also be considered to adapt the interface.
@@ B.1.1. Example system prompt
[No prose]
[Output only JSON]
Do not write normal text
You are a JSON generator which converts meeting agenda text into a more descriptive agenda
description. You always need to have an introduction phase at the beginning.
@@ B.1.2. Example user prompt
Please breakdown the following meeting agenda that someone has sent in email into meeting phases
that we would need to go through. We have a 60 minute meeting scheduled. Based on the given
 information, give a goal of the meeting (goal), as well as the explanation on why you chose the
 goal (exp). And also give the phase definition in a list (pi). Each phase definition should
 include: (1) Phase title (pt), (2) Phase description (pd) which should include a sub goal of
  a phase, (3) Behaviors to be encouraged (be), (4) Behaviors to be discouraged (bd), (5)
  priority (p), (6) amount of time allocation (t), (7) direction (d) (i.e., is it an iterative phase
   or directional phase). Please only respond in JSON with
each element needed as a key within a phase i.e., if we have two phases,
[{"pt":"xxx","pd":"xxx","be":["xxx","yyy"],
"bd":["xxx","yyy"],"p":"high","t":2,"d":"iterative"},
{"pt":"yyy","pd":"zzz","be":["ttt","kk"],
"bd":["lll","mmm"],"p":"low","t":8,"d":"directional"}].
So the overall JSON to export is {"goal":"xxx","pi":
[<phase definitions>],"exp":"xxx"}. Explanation should start by saying this goal is generated... or
 similar. Please use the full 60 minutes. Here is the meeting invitation:
(The email invitation would be attached.)
Refinement could be done using the same script, but by attaching the refinement scenario. This script can be
generated through code, and does not have to be detailed.
@@ B.2.1. Example system prompt
[No prose]
[Output only JSON]
Do not write normal text
You are a helpful assistant who creates screen layout that has appropriate apps that are most helpful
for users to complete the task successfully. Respond only in JSON following the format. Example
format:
[{"PhaseTitle":"xxx","timer":n,"programList":
[{"name":"yyy,"description":"zzz"},
{"name:"kk","description":"lll"}]},
{"PhaseTitle":"xxx","timer":n,"programList":
[{"name":"yyy,"description":"zzz"},
{"name:"kk","description":"lll"}]}].
Strictly follow this format. n is integer, and
programList.name should either be a name of a program in a program list given or a URL. programList.
description is where you should put a extremely concise reason why you chose that program. e.g.,
 Use this for presenting agenda; Use this for viewing relevant budget data
@@ B.2.2. Example user prompt
    I will give you the list of phases in a meeting in JSON format. Each phase in JSON is defined
    with several keywords.
    "pt" represents phase title, "pd" represents phase description, "be" represents behaviours to be
     encouraged, "bd" represents the behaviours to be
    discouraged, "p" represents priority (high, medium, low"), "t" represents recommended duration of
 time for the phase, and "d" represents directionality (directional i.e., cannot be returned,
 and should be preceeded by a certain phase or iterative i.e., can be transitioned to this whenever).
You need to generate what kind of programs are needed for helping goals of each phase (defined by
 the description) to be met the most efficiently. You can generate a list of 1-5 program name/URL,
  and the sequence of generation will affect where they are being placed, and size.
  Therefore, you need to be sensible about ordering so that important
 programs can be shown with the bigger sizes. The rule is as follows: If you have one program
 on the list, that would be full screen. If you have two programs, it would be one on the left half
 (first program on the list), and one on the right half. If you have three programs,
 the right-hand-side panel will split in half, creating two small panels at the top and the bottom.
  Four programs mean the left-hand-side panel will also be  split. Five programs mean two equally sized
   panel at the top, and three equally sized panel at the bottom.
   The ordering in the list will be used to place program to panels in a clockwise ordering (top
left panel is the first panel). Here is the list of programs available, and if the program that you
 want is not listed, please generate a URL for the program that you need instead of the program name.
Please feel free to give Bing Search URL with the search term filled, and generate at least one URL:
@@ B.3.1. Example system prompt for a specialized calculator generation
   You need to generate a HTML page, and only HTML+CSS
   +JavaScript based page as a response which allows me to calculate the total value to have the a
 list of features for a given scenario. Each feature needs to have an "include" button
     (green tick) or "exclude" button (red cross), and you need to calculate the total at the end,
      when the submit is clicked. You should not show any prices including itemised ones before this.
       You need to generate at least 30 features (features should be unique and descriptive. No Feature
        26 or something likethat), and incorporate that into the page, those are relevant for the scenario.
         Assign random prices for each feature. The HTML page needs to incorporate all the features that
          you generated embedded (i.e., no ... or "many more features here" etc.). No prose. No add more
           features here etc. You need to list all the features on the HTML.
@@ B.3.2. Example user prompt
   Designer, software engineer, hardware engineer, PM, marketing expert, and a researcher are
 gathering to hink about what features that a new headphone product that they release might have.
 It could be electronics feature such as active noise cancelling (95%) or it could be about
  materials (leather etc.).

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Designing Interactive Systems Conference; July 1–5, 2024; IT University of Copenhagen, Denmarkbooktitle: Designing Interactive Systems Conference (DIS ’24), July 1–5, 2024, IT University of Copenhagen, Denmarkdoi: 10.1145/3643834.3661507isbn: 979-8-4007-0583-0/24/07ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in interaction designccs: Human-centered computing Collaborative and social computing systems and tools
@@ 6.2.1. Linearity of meetings
Our findings suggest that both linear and non-linear meetings are common for our study participants, leading to mixed responses about surfacing the implicit phase structure in meetings. Some participants valued the structure that CoExplorer imposed on meetings by surfacing meeting goals and phases, helping to clarify attendee perceptions in advance of meetings, and supporting organizers in driving the meeting. This was perceived to improve meeting clarity and therefore productivity. However, some participants expressed concerns about the excessive structure imposed by CoExplorer, worrying that it may limit the natural spontaneity and flexibility of conversations. Considering these individual perceptions and preferences, it may be crucial for meeting support systems to capture the general consensus on the desired level of structure and adapt the workflow accordingly, rather than adopting a one-size-fits-all approach. Reconciling the variations in the group consensus and the individual preferences is a direction that needs further research. Systems could also help people decide on how much structure may be needed depending on their goals. It also points out the value in developing ways to adaptively adjust the level of structure during meetings. For instance, when people are moving through phases smoothly, the suggestion of structure could be more restrained, whereas if the meeting is being side-tracked, structure could be suggested in a stronger way by tracking the progress of the meeting through an intelligent system. Design implication 2: The design of intelligent meeting systems should consider ways to address the tension between structured, linear meetings enabling intentionality and clarity, and less structured, non-linear meetings, enabling flexibility. This includes capturing attendees’ preferences about structure as well as helping decide how much structure may be beneficial for their goals. Variations in need for structure during a given meeting should also be considered to adapt the interface.
@@ B.1.1. Example system prompt
{spverbatim}
[No prose] [Output only JSON] Do not write normal text You are a JSON generator which converts meeting agenda text into a more descriptive agenda description. You always need to have an introduction phase at the beginning.
@@ B.1.2. Example user prompt
{spverbatim}
Please breakdown the following meeting agenda that someone has sent in email into meeting phases that we would need to go through. We have a 60 minute meeting scheduled. Based on the given information, give a goal of the meeting (goal), as well as the explanation on why you chose the goal (exp). And also give the phase definition in a list (pi). Each phase definition should include: (1) Phase title (pt), (2) Phase description (pd) which should include a sub goal of a phase, (3) Behaviors to be encouraged (be), (4) Behaviors to be discouraged (bd), (5) priority (p), (6) amount of time allocation (t), (7) direction (d) (i.e., is it an iterative phase or directional phase). Please only respond in JSON with each element needed as a key within a phase i.e., if we have two phases, [”pt”:”xxx”,”pd”:”xxx”,”be”:[”xxx”,”yyy”], ”bd”:[”xxx”,”yyy”],”p”:”high”,”t”:2,”d”:”iterative”, ”pt”:”yyy”,”pd”:”zzz”,”be”:[”ttt”,”kk”], ”bd”:[”lll”,”mmm”],”p”:”low”,”t”:8,”d”:”directional”]. So the overall JSON to export is ”goal”:”xxx”,”pi”: [¡phase definitions¿],”exp”:”xxx”. Explanation should start by saying this goal is generated… or similar. Please use the full 60 minutes. Here is the meeting invitation: (The email invitation would be attached.)
Refinement could be done using the same script, but by attaching the refinement scenario. This script can be generated through code, and does not have to be detailed.
@@ B.2.1. Example system prompt
{spverbatim}
[No prose] [Output only JSON] Do not write normal text You are a helpful assistant who creates screen layout that has appropriate apps that are most helpful for users to complete the task successfully. Respond only in JSON following the format. Example format: [”PhaseTitle”:”xxx”,”timer”:n,”programList”: [”name”:”yyy,”description”:”zzz”, ”name:”kk”,”description”:”lll”], ”PhaseTitle”:”xxx”,”timer”:n,”programList”: [”name”:”yyy,”description”:”zzz”, ”name:”kk”,”description”:”lll”]]. Strictly follow this format. n is integer, and programList.name should either be a name of a program in a program list given or a URL. programList.description is where you should put a extremely concise reason why you chose that program. e.g., Use this for presenting agenda; Use this for viewing relevant budget data
@@ B.2.2. Example user prompt
{spverbatim}
I will give you the list of phases in a meeting in JSON format. Each phase in JSON is defined with several keywords. ”pt” represents phase title, ”pd” represents phase description, ”be” represents behaviours to be encouraged, ”bd” represents the behaviours to be discouraged, ”p” represents priority (high, medium, low”), ”t” represents recommended duration of time for the phase, and ”d” represents directionality (directional i.e., cannot be returned, and should be preceeded by a certain phase or iterative i.e., can be transitioned to this whenever). You need to generate what kind of programs are needed for helping goals of each phase (defined by the description) to be met the most efficiently. You can generate a list of 1-5 program name/URL, and the sequence of generation will affect where they are being placed, and size. Therefore, you need to be sensible about ordering so that important programs can be shown with the bigger sizes. The rule is as follows: If you have one program on the list, that would be full screen. If you have two programs, it would be one on the left half (first program on the list), and one on the right half. If you have three programs, the right-hand-side panel will split in half, creating two small panels at the top and the bottom. Four programs mean the left-hand-side panel will also be split. Five programs mean two equally sized panel at the top, and three equally sized panel at the bottom. The ordering in the list will be used to place program to panels in a clockwise ordering (top left panel is the first panel). Here is the list of programs available, and if the program that you want is not listed, please generate a URL for the program that you need instead of the program name. Please feel free to give Bing Search URL with the search term filled, and generate at least one URL:
@@ B.3.1. Example system prompt for a specialized calculator generation
{spverbatim}
You need to generate a HTML page, and only HTML+CSS +JavaScript based page as a response which allows me to calculate the total value to have the a list of features for a given scenario. Each feature needs to have an ”include” button (green tick) or ”exclude” button (red cross), and you need to calculate the total at the end, when the submit is clicked. You should not show any prices including itemised ones before this. You need to generate at least 30 features (features should be unique and descriptive. No Feature 26 or something like that), and incorporate that into the page, those are relevant for the scenario. Assign random prices for each feature. The HTML page needs to incorporate all the features that you generated embedded (i.e., no … or ”many more features here” etc.). No prose. No add more features here etc. You need to list all the features on the HTML.
@@ B.3.2. Example user prompt
{spverbatim}
Designer, software engineer, hardware engineer, PM, marketing expert, and a researcher are gathering to think about what features that a new headphone product that they release might have. It could be electronics feature such as active noise cancelling (95